The Eastside Nursery ordered 27 trees. Some of the trees were elms, costing $17 each. The remainder was maples at $11 each. The total cost of the trees was $375. Find the number of elms and the number of maples.
Answer by stanbon(75887)

The Eastside Nursery ordered 27 trees. Some of the trees were elms, costing $17 each. The remainder was maples at $11 each. The total cost of the trees was $375. Find the number of elms and the number of maples.
---------------------------------
Equations:
e + m = 27
17e+11m = 375
---
Multiply thru the 1st equation by 17:
17e + 17m = 17*27
17e + 11m = 375
----
Subtract and solve for "m":
6m = 84
m = 14 (# of maple trees purchased)
----
Solve for "e":
e + m = 27
e + 14 = 27
e = 13 (# of elm trees purchased)
=====================================
Cheers,
Stan H.
==========